# Magnetic field amplification during a turbulent collapse

### v2

@@ Abstract
The question of whether a dynamo can be triggered by gravitational collapse is of great interest, especially for the early Universe. Here, we employ supercomoving coordinates to study the magnetic field amplification from decaying turbulence during gravitational collapse. We perform three-dimensional simulations and show that for large magnetic Reynolds numbers there can be exponential growth of the comoving magnetic field with conformal time before the decay of turbulence impedes further amplification. The collapse dynamics only affects the nonlinear feedback from the Lorentz force, which diminishes more rapidly for shorter collapse times, allowing nearly kinematic continued growth. We confirm that helical turbulence is more efficient in driving dynamo action than nonhelical turbulence, but this difference decreases for larger collapse times. We also show that for nearly irrotational flows, dynamo amplification is still possible, but it is always associated with a growth of vorticity—even if it still remains very small. In nonmagnetic runs, the growth of vorticity is associated with viscosity and grows with the Mach number. In the presence of magnetic fields, vorticity emerges from the curl of the Lorentz force. During a limited time interval, an exponential growth of the comoving magnetic field with conformal time is interpreted as clear evidence of dynamo action.
@@ 1 Introduction
The hypothesis that dynamo action is ubiquitous in astrophysical plasmas was introduced in the 1950s, but it faced skepticism due to various anti-dynamo theorems (Cowling, 1933; Hide & Palmer, 1982). While initially the community focused on large-scale dynamos in the Sun (Parker, 1955; Steenbeck et al., 1966) and galaxies (Parker, 1971; Vainshtein & Ruzmaikin, 1971), the advance of powerful computers brought significant attention to small-scale dynamos at the scale of turbulence; see Meneguzzi et al. (1981) for the first simulations and Kazantsev (1968) for the underlying theory, as well as Kulsrud & Anderson (1992) for an independent and more detailed derivation. By now, it is clear that three-dimensional turbulence always leads to dynamo action when the plasma is sufficiently well conducting; see Brandenburg & Ntormousi (2023) for a recent review. This behavior implies that part of the kinetic energy in the turbulence is almost always converted into magnetic energy.
Collapse flows are particularly compelling for dynamo action. Since gravitational collapse provides a strong source of kinetic energy, it can enhance the magnetization of collapsing structures by sustaining or introducing turbulence in the flow. This mechanism is very relevant for galactic magnetism. Recently, there have been claims of strong ($\sim\mu$G or stronger) large-scale coherent galactic magnetic fields at redshifts up to 5.6 (Geach et al., 2023; Chen et al., 2024). Assuming only tiny primordial seeds magnetic fields, there might not be enough time for a high redshift galaxy to build strong enough magnetic fields through mean-field dynamo action. An early amplification of a tiny initial seed through a small-scale dynamo (Beck et al., 1994), especially during the gravitational collapse of the initial halo, could alleviate this problem.
@@
Despite its relevance to various astrophysical environments, gravitational collapse dynamos have not yet been convincingly demonstrated. The main reason is that characterizing dynamos in unsteady flows is inherently challenging. For steady flows, we can always formulate an eigenvalue problem, provided the magnetic field is still weak and unaffected by the feedback from the Lorentz force, which affects the flow amplitude. It is even possible to prove that there is no eigenfunction with a non-vanishing eigenvalue when the magnetic diffusivity is strictly zero (Moffatt & Proctor, 1985). Unsteady flows present a significant complication because, in that situation, the kinematic growth or decay of the magnetic field is no longer exponential. The problem becomes approachable if the flow is statistically steady, i.e., the level of turbulence remains constant over time. In such cases, the energy spectrum grows at all wavenumbers at the same rate (Subramanian & Brandenburg, 2014). This behavior is suggestive of the existence of an eigenfunction of the type discussed by Kazantsev (1968). However, many astrophysical flows, such as gravitational collapse, are not even statistically steady. Dynamo research in these cases is still in its infancy.
@@
Irshad P et al. (2025) found super-exponential growth of the magnetic field as a result of the increasing turnover rate and saturation field strengths over the expectations from flux freezing. They applied a solenoidal forcing function with and without kinetic helicity. The present work aims to study decaying turbulence during gravitational collapse by employing supercomoving coordinates and allowing not only for cases without initial kinetic helicity but also cases with or without initial vorticity, i.e., acoustic turbulence.
@@ 2.1 Supercomoving coordinates
where $t$ is the conformal time, $s$ is a free-fall parameter, which is related to the free-fall time $t_{\rm ff}=\pi/2s$. The physical time $t_{\mathrm{ph}}$ is then given by
@@ 2.3 Initial conditions and parameters
We construct our initial velocity in Fourier space (indicated by a tilde) as $\tilde{\bm{u}}(\bm{k})=\bm{{\mathsf{M}}}(\bm{k})\bm{S}(\bm{k})$. Here,
@@
Run $\mathcal{S}$ $\sigma$ $\zeta$ $k_{0}/k_{1}$ $\mbox{\rm Ma}_{0}$ $\mbox{\rm Re}_{\rm M}$ (Re) $\mathcal{B}_{\mathrm{ini}}$ $\Delta\ln\mathcal{B}$ $\lambda/u_{0}k_{0}$ $\Delta\ln(k_{\omega}/k_{0})$ $(k_{\omega}/k_{0})_{\max}$ resol. 1 0.1 1 0 10 0.18 1840 $2.3\times 10^{-8}$ 8.33 0.52 0.39 7.09 $512^{3}$ 2 0.1 1 0 10 0.18 1840 $2.3\times 10^{-5}$ 6.62 0.52 0.39 7.09 $512^{3}$ 3 0.1 1 0 10 0.18 1840 $2.3\times 10^{-2}$ 1.88 1.00 0.31 6.46 $512^{3}$ 4 0.6 1 0 10 0.18 1840 $2.3\times 10^{-2}$ 2.21 1.03 0.22 5.93 $512^{3}$ 5 2.8 1 0 10 0.18 1840 $2.3\times 10^{-2}$ 3.56 1.03 0.30 6.43 $512^{3}$ 6 11 1 0 10 0.18 1840 $2.3\times 10^{-2}$ 4.77 1.03 0.36 6.82 $512^{3}$ 7 56 1 0 10 0.18 1840 $2.3\times 10^{-2}$ 5.96 1.03 0.39 7.04 $512^{3}$ 8 0.2 0 0 10 0.13 1300 $3.3\times 10^{-8}$ 4.27 0.37 0.33 6.97 $512^{3}$ 9 0.2 0 0 10 0.13 1300 $3.3\times 10^{-5}$ 4.22 0.37 0.33 6.97 $512^{3}$ 10 0.2 0 0 10 0.13 1300 $3.3\times 10^{-2}$ 1.49 0.97 0.14 5.70 $512^{3}$ 11 0.8 0 0 10 0.13 1300 $3.3\times 10^{-2}$ 1.92 0.97 0.17 5.91 $512^{3}$ 12 3.8 0 0 10 0.13 1300 $3.3\times 10^{-2}$ 3.03 0.98 0.29 6.66 $512^{3}$ 13 15 0 0 10 0.13 1300 $3.3\times 10^{-2}$ 3.75 0.98 0.33 6.92 $512^{3}$ 14 77 0 0 10 0.13 1300 $3.3\times 10^{-2}$ 4.12 0.98 0.33 6.97 $512^{3}$ 15 0.2 0 0.10 10 0.12 1170 $3.6\times 10^{-2}$ 1.41 0.34 0.11 5.50 $512^{3}$ 16 0.2 0 0.50 10 0.08 800 $5.4\times 10^{-2}$ 1.04 0.25 0.00 4.00 $512^{3}$ 17 0.2 0 0.90 10 0.08 840 $5.1\times 10^{-2}$ 0.31 0.04 0.25 0.94 $512^{3}$ 18 0.2 0 0.95 10 0.09 880 $4.9\times 10^{-2}$ 0.05 0.003 0.28 0.47 $512^{3}$ 19 0.2 0 0.96 10 0.09 880 $4.8\times 10^{-2}$ 0.02 0.001 0.26 0.38 $512^{3}$ 20 0.2 0 0.97 10 0.09 890 $4.8\times 10^{-2}$ 0.21 0.29 $512^{3}$ 21 0.2 0 0.98 10 0.09 900 $4.7\times 10^{-2}$ 0.13 0.20 $512^{3}$ 22 0.2 0 0.99 10 0.09 910 $4.7\times 10^{-2}$ 0.20 0.16 $512^{3}$ 23 0.2 0 1 10 0.09 920 $4.6\times 10^{-2}$ 0.30 0.14 $512^{3}$ 24 0.1 0 1 20 0.09 920 0.01 0.07 $1024^{3}$ 25 0.2 0 1 10 0.09 930 0.03 0.05 $1024^{3}$ 26 0.4 0 1 5 0.09 940 0.38 0.04 $1024^{3}$ 27 1.0 0 1 2 0.10 950 1.27 0.03 $1024^{3}$ 28 0.5 0 0.95 10 0.04 220 0.09 0.23 $512^{3}$ 29 0.1 0 0.95 10 0.18 890 0.31 0.71 $1024^{3}$ 30 0.1 0 0.95 10 0.27 1330 0.43 1.00 $1024^{3}$ 31 0.1 0 0.95 10 0.36 1780 0.51 1.31 $1024^{3}$ 32 0.2 0 0.96 10 0.09 900 $4.9\times 10^{-2}$ 0.02 0.001 0.17 0.38 $1024^{3}$ 33 0.2 0 0.96 10 0.09 1800 $4.9\times 10^{-2}$ 0.12 0.004 0.28 0.53 $1024^{3}$ 34 0.2 0 0.96 10 0.09 4500 $4.9\times 10^{-2}$ 0.51 0.008 0.53 0.79 $1024^{3}$ 35 0.2 0 1 10 0.09 1870 $9.4\times 10^{-3}$ 0.03 0.07 $1024^{3}$ 36 0.2 0 1 10 0.09 1870 $2.4\times 10^{-2}$ 0.17 0.09 $1024^{3}$ 37 0.2 0 1 10 0.09 1870 $4.7\times 10^{-2}$ 0.34 0.21 $1024^{3}$ 38 0.2 0 1 10 0.09 1870 $9.4\times 10^{-2}$ 0.25 0.48 $1024^{3}$ 39 0.4 1 0 10 0.19 190 $2.3\times 10^{-17}$ 8.32 0.42 0.01 4.29 $2048^{3}$
@@ 3 Results
3.1 Growth vs physical and conformal time
@@
In physical time, the magnetic field shows a steep increase just toward the end of the collapse; see Figure 2. Interestingly, the runs with large values of $\mathcal{S}$, which produce the strongest comoving magnetic fields, now yield the weakest physical fields when comparing the runs at the same fractional collapse time. This is because for the runs with large values of $\mathcal{S}$, the free-fall time is short, so the fractional times are larger, which effectively interchanges the order of the curves. This is demonstrated in the inset of Figure 2, where we show the same data, but now with time in units of the initial turnover time.
In Figure 2, we have also indicated the times where the initial exponential growth of the comoving magnetic field with conformal time terminates. For $\mathcal{S}=0.1$ and 0.6, $\mathcal{B}_{\rm rms}/a^{2}$ has hardly increased by an order of magnitude. In particular, the growth of $\mathcal{B}_{\rm rms}/a^{2}$ versus physical time is not super-exponential, as found by Irshad P et al. (2025). The reason for our subexponential growth for $\mathcal{S}\ll 1$ is that the rms velocity decreases significantly due to turbulent diffusion leading to a smaller growth rate which then counters the effect of collapse. Only for larger values of $\mathcal{S}$ is the growth super-exponential in physical coordinates, and exponential in comoving coordinates. For $\mathcal{S}\geq 2.8$, the times when exponential growth in comoving coordinates terminates are outside the plot range of Figure 2.
Figure 2: Same as Figure 1, but in physical units. Time is here normalized by the free-fall time. The black and blue dots on the black and blue curves denote the time until which the growth in Figure 1 was still approximately exponential. The inset shows the same, but now time is normalized by the initial turnover time. Runs 3–7 and Runs 10–14.
@@ 3.2 Effect of the Lorentz force
[…]
@@
Figure 4: (a) Similar to Figure 1, but now $a^{1/2}\mathcal{B}_{\rm rms}$ (thicker lines) and the instantaneous rms velocity (thinner lines) are plotted. The order of the colors is the same as before, with black being for $s/c_{\rm s}k_{1}=0.2$ and red for $s/c_{\rm s}k_{1}=100$ and solid (dashed) lines refer to helical (nonhelical) initial flows, for which $\mathcal{S}$ varies in the range 0.1–56 (0.2–77). (b) Evolution separately for $a^{1/2}$ (dashed-dotted lines) and $\mathcal{B}_{\rm rms}$ (solid lines), again with the same colors as before. Runs 3–7 and Runs 10–14.
We see that Equation (12) is well obeyed for all runs. The largest values of $a^{1/2}\mathcal{B}_{\rm rms}$ are obtained for the runs with small values of $\mathcal{S}$. The effect of kinetic helicity is here surprisingly weak and the values of $a^{1/2}\mathcal{B}_{\rm rms}$ are only slightly smaller for the nonhelical runs than for the helical ones. For larger values of $\mathcal{S}$, on the other hand, the differences between helical and nonhelical runs are much larger and we see that the decay of $a^{1/2}$ is well overcompensated by the growth of $\mathcal{B}_{\rm rms}$ so that the product $a^{1/2}\mathcal{B}_{\rm rms}$ still shows a strong increase later in the evolution; see Figure 4(b), where we plot separately the evolutions of $a^{1/2}$ and $\mathcal{B}_{\rm rms}$.

### v3

@@ Abstract
The question of whether a dynamo can be triggered by gravitational collapse is of great interest, especially for the early Universe. Here, we employ supercomoving coordinates to study the magnetic field amplification from decaying turbulence during gravitational collapse. We perform three-dimensional simulations and show that for large magnetic Reynolds numbers, there can be exponential growth of the comoving magnetic field with conformal time before the decay of turbulence impedes further amplification. The collapse dynamics only affect the nonlinear feedback from the Lorentz force, which diminishes more rapidly for shorter collapse times, allowing nearly kinematic continued growth. We confirm that helical turbulence is more efficient in driving dynamo action than nonhelical turbulence, but this difference decreases for larger collapse times. We also show that for nearly irrotational flows, dynamo amplification is still possible, but it is always associated with a growth of vorticity—even if it still remains very small. In nonmagnetic runs, the growth of vorticity is associated with viscosity and grows with the Mach number. In the presence of magnetic fields, vorticity emerges from the curl of the Lorentz force. During a limited time interval, an exponential growth of the comoving magnetic field with conformal time is interpreted as clear evidence of dynamo action.
@@ 1 Introduction
The hypothesis that dynamo action is ubiquitous in astrophysical plasmas was introduced in the 1950s, but it faced skepticism due to various antidynamo theorems (Cowling, 1933; Hide & Palmer, 1982). While initially the community focused on large-scale dynamos in the Sun (Parker, 1955; Steenbeck et al., 1966) and galaxies (Parker, 1971; Vainshtein & Ruzmaikin, 1971), the advance of powerful computers brought significant attention to small-scale dynamos at the scale of turbulence; see Meneguzzi et al. (1981) for the first simulations and Kazantsev (1968) for the underlying theory, as well as Kulsrud & Anderson (1992) for an independent and more detailed derivation. By now, it is clear that three-dimensional turbulence always leads to dynamo action when the plasma is sufficiently well conducting; see Brandenburg & Ntormousi (2023) for a recent review. This behavior implies that part of the kinetic energy in the turbulence is almost always converted into magnetic energy.
[…]
@@
Despite its relevance to various astrophysical environments, gravitational collapse dynamos have not yet been convincingly demonstrated. The main reason is that characterizing dynamos in unsteady flows is inherently challenging. For steady flows, we can always formulate an eigenvalue problem, provided the magnetic field is still weak and unaffected by the feedback from the Lorentz force, which affects the flow amplitude. It is even possible to prove that there is no eigenfunction with a nonvanishing eigenvalue when the magnetic diffusivity is strictly zero (Moffatt & Proctor, 1985). Unsteady flows present a significant complication because, in that situation, the kinematic growth or decay of the magnetic field is no longer exponential. The problem becomes approachable if the flow is statistically steady, i.e., the level of turbulence remains constant over time. In such cases, the energy spectrum grows at all wavenumbers at the same rate (Subramanian & Brandenburg, 2014). This behavior is suggestive of the existence of an eigenfunction of the type discussed by Kazantsev (1968). However, many astrophysical flows, such as gravitational collapse, are not even statistically steady. Dynamo research in these cases is still in its infancy.
@@
Irshad P et al. (2025) found superexponential growth of the magnetic field as a result of the increasing turnover rate and saturation field strengths over the expectations from flux freezing. They applied a solenoidal forcing function with and without kinetic helicity. The present work aims to study decaying turbulence during gravitational collapse by employing supercomoving coordinates and allowing not only for cases without initial kinetic helicity but also cases with or without initial vorticity, i.e., acoustic turbulence.
@@ 2.1 Supercomoving coordinates
where $t$ is the conformal time, $s$ is a freefall parameter, which is related to the freefall time $t_{\rm ff}=\pi/2s$. The physical time $t_{\mathrm{ph}}$ is then given by
@@ 2.3 Initial conditions and parameters
We construct our initial velocity in Fourier space (indicated by a tilde) as $\tilde{\bm{u}}(\bm{k})=\bm{{\mathsf{M}}}(\bm{k})\bm{S}(\bm{k})$. Here
@@
Run $\mathcal{S}$ $\sigma$ $\zeta$ $k_{0}/k_{1}$ $\mbox{\rm Ma}_{0}$ $\mbox{\rm Re}_{\rm M}$ (Re) $\mathcal{B}_{\mathrm{ini}}$ $\Delta\ln\mathcal{B}$ $\lambda/u_{0}k_{0}$ $\Delta\ln(k_{\omega}/k_{0})$ $(k_{\omega}/k_{0})_{\max}$ resol. 1 0.1 1 0 10 0.18 1840 $2.3\times 10^{-8}$ 8.33 0.52 0.39 7.09 $512^{3}$ 2 0.1 1 0 10 0.18 1840 $2.3\times 10^{-5}$ 6.62 0.52 0.39 7.09 $512^{3}$ 3 0.1 1 0 10 0.18 1840 $2.3\times 10^{-2}$ 1.88 1.00 0.31 6.46 $512^{3}$ 4 0.6 1 0 10 0.18 1840 $3\times 10^{-2} 2.2$1 1.03 0.22 5.93 $512^{3}$ 5 2.8 1 0 10 0.18 1840 $2.3\times 10^{-2}$ 3.56 1.03 0.30 6.43 $512^{3}$ 6 11 1 0 10 0.18 1840 $2.3\times 10^{-2}$ 4.77 1.03 0.36 6.82 $512^{3}$ 7 56 1 0 10 0.18 1840 $2.3\times 10^{-2}$ 5.96 1.03 0.39 7.04 $512^{3}$ 8 0.2 0 0 10 0.13 1300 $3.3\times 10^{-8}$ 4.27 0.37 0.33 6.97 $512^{3}$ 9 0.2 0 0 10 0.13 1300 $3.3\times 10^{-5}$ 4.22 0.37 0.33 6.97 $512^{3}$ 10 0.2 0 0 10 0.13 1300 $3.3\times 10^{-2}$ 1.49 0.97 0.14 5.70 $512^{3}$ 11 0.8 0 0 10 0.13 1300 $3.3\times 10^{-2}$ 1.92 0.97 0.17 5.91 $512^{3}$ 12 3.8 0 0 10 0.13 1300 $3.3\times 10^{-2}$ 3.03 0.98 0.29 6.66 $512^{3}$ 13 15 0 0 10 0.13 1300 $3.3\times 10^{-2}$ 3.75 0.98 0.33 6.92 $512^{3}$ 14 77 0 0 10 0.13 1300 $3.3\times 10^{-2}$ 4.12 0.98 0.33 6.97 $512^{3}$ 15 0.2 0 0.10 10 0.12 1170 $3.6\times 10^{-2}$ 1.41 0.34 0.11 5.50 $512^{3}$ 16 0.2 0 0.50 10 0.08 800 $5.4\times 10^{-2}$ 1.04 0.25 0.00 4.00 $512^{3}$ 17 0.2 0 0.90 10 0.08 840 $5.1\times 10^{-2}$ 0.31 0.04 0.25 0.94 $512^{3}$ 18 0.2 0 0.95 10 0.09 880 $4.9\times 10^{-2}$ 0.05 0.003 0.28 0.47 $512^{3}$ 19 0.2 0 0.96 10 0.09 880 $4.8\times 10^{-2}$ 0.02 0.001 0.26 0.38 $512^{3}$ 20 0.2 0 0.97 10 0.09 890 $4.8\times 10^{-2}$ 0.21 0.29 $512^{3}$ 21 0.2 0 0.98 10 0.09 900 $4.7\times 10^{-2}$ 0.13 0.20 $512^{3}$ 22 0.2 0 0.99 10 0.09 910 $4.7\times 10^{-2}$ 0.20 0.16 $512^{3}$ 23 0.2 0 1 10 0.09 920 $4.6\times 10^{-2}$ 0.30 0.14 $512^{3}$ 24 0.1 0 1 20 0.09 920 0.01 0.07 $1024^{3}$ 25 0.2 0 1 10 0.09 930 0.03 0.05 $1024^{3}$ 26 0.4 0 1 5 0.09 940 0.38 0.04 $1024^{3}$ 27 1.0 0 1 2 0.10 950 1.27 0.03 $1024^{3}$ 28 0.5 0 0.95 10 0.04 220 0.09 0.23 $512^{3}$ 29 0.1 0 0.95 10 0.18 890 0.31 0.71 $1024^{3}$ 30 0.1 0 0.95 10 0.27 1330 0.43 1.00 $1024^{3}$ 31 0.1 0 0.95 10 0.36 1780 0.51 1.31 $1024^{3}$ 32 0.2 0 0.96 10 0.09 900 $4.9\times 10^{-2}$ 0.02 0.001 0.17 0.38 $1024^{3}$ 33 0.2 0 0.96 10 0.09 1800 $4.9\times 10^{-2}$ 0.12 0.004 0.28 0.53 $1024^{3}$ 34 0.2 0 0.96 10 0.09 4500 $4.9\times 10^{-2}$ 0.51 0.008 0.53 0.79 $1024^{3}$ 35 0.2 0 1 10 0.09 1870 $9.4\times 10^{-3}$ 0.03 0.07 $1024^{3}$ 36 0.2 0 1 10 0.09 1870 $2.4\times 10^{-2}$ 0.17 0.09 $1024^{3}$ 37 0.2 0 1 10 0.09 1870 $4.7\times 10^{-2}$ 0.34 0.21 $1024^{3}$ 38 0.2 0 1 10 0.09 1870 $9.4\times 10^{-2}$ 0.25 0.48 $1024^{3}$ 39 0.4 1 0 10 0.19 190 $2.3\times 10^{-17}$ 8.32 0.42 0.01 4.29 $2048^{3}$
@@ 3 Results
3.1 Growth versus physical and conformal time
@@
In physical time, the magnetic field shows a steep increase just toward the end of the collapse; see Figure 2. Interestingly, the runs with large values of $\mathcal{S}$, which produce the strongest comoving magnetic fields, now yield the weakest physical fields when comparing the runs at the same fractional collapse time. This is because for the runs with large values of $\mathcal{S}$, the freefall time is short, so the fractional times are larger, which effectively interchanges the order of the curves. This is demonstrated in the inset of Figure 2, where we show the same data, but now with time in units of the initial turnover time.
In Figure 2, we have also indicated the times where the initial exponential growth of the comoving magnetic field with conformal time terminates. For $\mathcal{S}=0.1$ and 0.6, $\mathcal{B}_{\rm rms}/a^{2}$ has hardly increased by an order of magnitude. In particular, the growth of $\mathcal{B}_{\rm rms}/a^{2}$ versus physical time is not superexponential, as found by Irshad P et al. (2025). The reason for our subexponential growth for $\mathcal{S}\ll 1$ is that the rms velocity decreases significantly due to turbulent diffusion, leading to a smaller growth rate, which then counters the effect of collapse. Only for larger values of $\mathcal{S}$ is the growth superexponential in physical coordinates, and exponential in comoving coordinates. For $\mathcal{S}\geq 2.8$, the times when exponential growth in comoving coordinates terminates are outside the plot range of Figure 2.
Figure 2: Same as Figure 1, but in physical units. Time is here normalized by the freefall time. The black and blue dots on the black and blue curves denote the time until which the growth in Figure 1 was still approximately exponential. The inset shows the same, but now time is normalized by the initial turnover time. Runs 3–7 and Runs 10–14.
@@ 3.2 Effect of the Lorentz force
$a^{1/2}B_{\rm rms}\mathrel{\mathchoice{\vbox{\offinterlineskip\halign{\hfil$\displaystyle#$\hfil\cr<\cr\sim\cr}}}{\vbox{\offinterlineskip\halign{\hfil$\textstyle#$\hfil\cr<\cr\sim\cr}}}{\vbox{\offinterlineskip\halign{\hfil$\scriptstyle#$\hfil\cr<\cr\sim\cr}}}{\vbox{\offinterlineskip\halign{\hfil$\scriptscriptstyle#$\hfil\cr<\cr\sim\cr}}}}u_{\rm rms}\sqrt{\mu_{0}\rho_{0}}.$ (12)
@@
Figure 4: (a) Similar to Figure 1, but now $a^{1/2}\mathcal{B}_{\rm rms}$ (thicker lines) and the instantaneous rms velocity (thinner lines) are plotted. The order of the colors is the same as before, with black being for $s/c_{\rm s}k_{1}=0.2$ and red for $s/c_{\rm s}k_{1}=100$ and solid (dashed) lines refer to helical (nonhelical) initial flows, for which $\mathcal{S}$ varies in the range 0.1–56 (0.2–77). (b) Evolution separately for $a^{1/2}$ (dashed–dotted lines) and $\mathcal{B}_{\rm rms}$ (solid lines), again with the same colors as before. Runs 3–7 and Runs 10–14.
We see that Equation (12) is well obeyed for all runs. The largest values of $a^{1/2}\mathcal{B}_{\rm rms}$ are obtained for the runs with small values of $\mathcal{S}$. The effect of kinetic helicity here is surprisingly weak, and the values of $a^{1/2}\mathcal{B}_{\rm rms}$ are only slightly smaller for the nonhelical runs than for the helical ones. For larger values of $\mathcal{S}$, on the other hand, the differences between helical and nonhelical runs are much larger and we see that the decay of $a^{1/2}$ is well overcompensated by the growth of $\mathcal{B}_{\rm rms}$ so that the product $a^{1/2}\mathcal{B}_{\rm rms}$ still shows a strong increase later in the evolution; see Figure 4(b), where we plot separately the evolutions of $a^{1/2}$ and $\mathcal{B}_{\rm rms}$.